Fanelli Corporation, a merchandising company, reported the following results for July:

Number of units sold 5,800

Selling price per unit $ 590

Unit cost of goods sold $ 410

Variable selling expense per unit $ 66

Total fixed selling expense $ 125,300

Variable administrative expense per unit $ 22

Total fixed administrative expense $ 207,200

Cost of goods sold is a variable cost in this company.

Required:

a. Prepare a traditional format income statement for July.

b. Prepare a contribution format income statement for July.
Answer #1

Answers

  • All working forms part of the answer
  • Requirement ‘a’

Traditional Income Statement

A = 5800 units x $ 590

Sales Revenue

$3,422,000

B = 5800 units x $ 410

Cost of Goods Sold

$2,378,000

C = A - B

Gross Profits

$1,044,000

Operating expenses:

Selling expenses--

D = 5800 units x $ 66

     Variable

$382,800

E

     Fixed

$125,300

Administrative expenses--

F = 5800 units x $ 22

     Variable

$127,600

G

     Fixed

$207,200

H = D+E+F+G

Total Operating expenses

$842,900

I = C - H

Net Operating income

$201,100

  • Requirement ‘b’

Contribution Format Income Statement

A = 5800 units x $ 590

Sales Revenue

$3,422,000

Variable costs:

B = 5800 units x $ 410

Cost of Goods Sold

$2,378,000

C = 5800 units x $ 66

Variable selling expenses

$382,800

D = 5800 units x $ 22

Variable administrative expenses

$127,600

E = B+C+D

Total variable cost

$2,888,400

F = A - E

Contribution margin

$533,600

Fixed Costs:

G

Fixed selling expense

$125,300

H

Fixed administrative expense

$207,200

I = G+H

Total Fixed Cost

$332,500

J = F - I

Net Operating Income

$201,100